# The Role of Integrity Monitoring in Connected and Automated Vehicles: Current State-of-Practice and Future Directions

### v1

@@ Abstract
Connected and Automated Vehicle (CAV) research has gained traction in the last decade due to significant advancements in perception, navigation, communication, and control functions. Accurate and reliable position information is needed to meet the requirements of CAV applications, especially when safety is concerned. With the advent of various perception sensors (e.g. camera, LiDAR, etc.), the vehicular positioning system has improved both in accuracy and robustness. Vehicle-to-Vehicle (V2V) and Vehicle-to-Infrastructure (V2I) based cooperative positioning can improve the accuracy of the position estimates, but the integrity risks involved in multi-sensor fusion in a cooperative environment have not yet been fully explored. This paper reviews existing research in the field of positioning Integrity Monitoring (IM) and identifies various research gaps. Particular attention has been placed on identifying research that highlights cooperative IM methods. This analysis helps pave the way for the development of new IM frameworks for cooperative positioning solutions in the future.
@@ I Introduction
Connected and Automated Vehicle (CAV) research is crucial to ensure safety, smooth traffic operation, and minimize energy consumption and carbon emissions. With the advent of sophisticated sensor platforms, computation hardware, and communication capability, it is now possible to deploy intelligent CAV applications through cooperation between various road agents. CAV applications can be broadly classified into three major use categories: safety, mobility, and environmental. Mobility and environmental applications such as eco-routing, eco-approach and departure (EAD) typically require coarse positioning accuracy (5-10 m) and/or lane level positioning accuracy (<1absent1<1< 1 m), while safety-critical applications such as collision avoidance, autonomous intersection management, etc., generally require a where-in lane level accuracy (<<< 0.2 m). Nigel et al. provided a qualitative analysis on various CAV applications and their positioning requirements in  [1]. The positioning accuracy required for various CAV applications can be achieved through onboard positioning sensors, such as the Global Navigation Satellite System (GNSS) receivers and Inertial Measurement Units (IMU), as well as perception sensors like cameras and LiDAR. Additionally, external sensor data, including information from roadside cameras and LiDAR, can be shared via wireless communication technologies such as cellular and/or Dedicated Short Range Communication (DSRC).
Further, in a mixed traffic environment, vehicles can be classified as Human-driven Vehicles (HDV), Automated Vehicles (AV), Connected Vehicles (CV), or Connected and Automated Vehicles (CAV). Various levels of autonomy (SAE Level 0-5) [2] and connectivity (fully or partially connected) also exist for CAVs. Mixed traffic also consists of vulnerable road users (VRUs) who are often not given much attention in CAV research. They are assumed to be non-connected and are tracked by vehicular and infrastructure sensors. This is not entirely true as many VRUs typically possess smart devices such as smartphones and smartwatches which have built-in GNSS/IMU capabilities with cellular connectivity. These features can facilitate VRU cooperation with the rest of the traffic improving their safety. Although mixed traffic encompasses a huge variety of road agents, only vehicles are considered in this review.
Various types of sensors are used to estimate the state of a vehicle and perceive the environment around it. Providing accurate vehicle state information while maintaining a lower risk of misinformation is instrumental in smooth navigation and control operations. Several integrity monitoring (IM) techniques have been used to monitor key parameters such as accuracy, availability, continuity, and integrity risk for positioning systems. The historical development of IM methods revolves around the Receiver Autonomous Integrity Monitoring (RAIM) method [3] development in the 1980s as a way to evaluate the trustworthiness of GPS signals used in aviation and marine navigation. Recently, more attention has been received in cooperative driving automation (CDA) research, as the existence of mixed traffic, varied sensing, and compute capabilities makes it challenging to implement and evaluate the integrity for real-time safety-critical CAV applications.
The type, quantity, placement, and measurement uncertainty of sensors play an important role in determining the navigation performance of a vehicle. Onboard positioning systems utilizing GNSS-IMU modules have proven to be a good source for providing absolute position information. In deep urban canyons, additional perception sensors such as a camera or LiDAR can augment these GNSS-based positioning systems, when signals are lost. Perception sensors are expensive, but they play an important role for autonomous driving functions in vehicles. Legacy vehicles or HDVs may not possess the necessary sensors for autonomous functionalities, but they can still reap the benefits of a connected ecosystem by sharing information with surrounding vehicles and infrastructure via vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communications. Today, many vehicle owners use smartphones for plug-in navigation applications such as Android Auto and Apple Carplay [4, 5]. With cellular connectivity, HDVs are loosely connected to the cellular ecosystem and can thus share information over the network, making them a partially-connected vehicle (P-CV). The use of smartphone-based plug-in technology has primarily been explored for long-term planning, such as route planning. However, its potential for reactive or short-term planning, particularly in safety-critical scenarios, remains largely untapped.
AVs and CAVs have access to a large amount of sensor data, but processing them in real time can be challenging. AVs have access to the onboard sensor data through reliable local wired connections. This allows an AV to run algorithms on a centralized on-board computation unit. However, in a cooperative or collaborative environment, the sensor information from multiple vehicles/infrastructure nodes can be processed either in a centralized or decentralized manner. This depends on the computational capabilities of the individual nodes (vehicle, infrastructure), communication capacity, and the end requirements of the vehicle.
In a mixed-traffic scenario, there exist various types of positioning solutions, vehicles, sensor modalities, and applications. It is unlikely that two vehicles having the same set of sensors and computation hardware will produce similar navigation performance. Hence, it is important to study and analyze the integrity of positioning systems under various conditions.
@@
The different sections of this paper are organized as follows. Section II discusses various uncertainties considered in past CAV research, i.e. model uncertainty, position uncertainty, etc. Integrity monitoring methods for vehicles via standalone and cooperative operations are presented in Section III. Section IV discusses key automotive safety standards in practice. Section VI explores existing research gaps and highlights potential opportunities for future work. The paper concludes with a comprehensive conclusion, synthesizing the key findings and implications of the research in Section VII.
@@ II Background: Consideration of Positioning Uncertainties in CAV Applications
Figure 1: Fundamental operations of vehicle platooning.
@@
Figure 2: Research Gaps in CAV applications.
@@ III Positioning Integrity Monitoring
The integrity monitoring (IM) framework consists of a set of requirements for the navigation system that must be met to ensure safe operation. These requirements are subject to the positioning system used by the vehicle and the application it intends to perform. These requirements are referred to as Required Navigation Performance (RNP). The performance of the vehicle positioning and navigation system is quantified using four properties defined by the RNP framework - Accuracy, Availability, Continuity, and Integrity. These properties are quantified by four parameters [25, 26]:
@@
Figure 3: Stanford-ESA integrity diagram to monitor positioning integrity [26].
@@
The Stanford-ESA (Error Separation Algorithm) integrity diagram [26] is a graphical tool used to illustrate and assess the integrity of GNSS-based navigation systems, particularly in aviation. Although this plot was developed for GNSS positioning systems, it can be adapted and used for other navigation solutions. The PL estimated from standard deviations in position estimates and multipliers derived from desired confidence levels is plotted against the true position error over time across different scenarios. Different operating zones in the Stanford plot are discussed in detail in [25] and [27]. Integrity risk is determined by the percentage of (PL, PE) pair in the Hazardously Misleading Information (HMI) region without raising an alert, thus compromising safety of the vehicle.
III-A In-vehicle (Standalone) Integrity Monitoring

### v2

@@ Abstract
Positioning integrity refers to the trust in the performance of a navigation system. Accurate and reliable position information is needed to meet the requirements of Connected and Automated Vehicle (CAV) applications, particularly in safety-critical scenarios. Receiver Autonomous Integrity Monitoring (RAIM) and its variants have been widely studied for Global Navigation Satellite System (GNSS)-based vehicle positioning, often fused with kinematic (e.g., Odometry) and perception sensors (e.g., camera). However, integrity monitoring (IM) for cooperative positioning solutions leveraging Vehicle-to-Everything (V2X) communication has received comparatively limited attention. This paper reviews existing research in the field of positioning IM and identifies various research gaps. Particular attention has been placed on identifying research that highlights cooperative IM methods. It also examines key automotive safety standards and public V2X datasets to map current research priorities and uncover critical gaps. Finally, the paper outlines promising future directions, highlighting research topics aimed at advancing and benchmarking positioning integrity.
@@ I Introduction
Connected and Automated Vehicle (CAV) research is crucial to ensure safety, smooth traffic operation, and minimize energy consumption and carbon emissions [1]. With the advent of sophisticated sensor platforms, computation hardware, and communication capability, it is now possible to deploy intelligent CAV applications through cooperation between various road agents. CAV applications can be broadly classified into three major use categories: safety, mobility, and environmental. Mobility and environmental applications such as eco-routing, eco-approach and departure (EAD) typically require coarse positioning accuracy (5-10 m) and/or lane level positioning accuracy ($<1$ m), while safety-critical applications such as collision avoidance, autonomous intersection management, etc., generally require a where-in lane level accuracy ($<$ 0.2 m). Nigel et al. provided a qualitative analysis on various CAV applications and their positioning requirements in  [2]. The positioning accuracy required for various CAV applications can be achieved through onboard positioning sensors, such as the Global Navigation Satellite System (GNSS) receivers and Inertial Measurement Units (IMU), as well as perception sensors like cameras and LiDAR. Additionally, Roadside camera and LiDAR data can be broadcast over cellular or DSRC links and incorporated via relative-ranging techniques to enhance cooperative positioning [3].
A CAV application’s feasibility is dependent on the positioning performance of a specific sensor platform (standalone positioning) in the vehicle and/or its ability to use the information shared by neighboring road agents (cooperative positioning). Standalone positioning techniques include GNSS-based positioning [4], which is often fused with other onboard kinematic sensors such as wheel odometry [5], Inertial Measurement Units (IMU) [6], and perception sensors such as radars, cameras, and LiDARs [7]. On the other hand, cooperative positioning relies on relative ranging techniques and cooperative feature association techniques for vehicle localization.
The positioning performance of a navigation system is characterized by four key parameters: accuracy, availability, continuity, and integrity risk [8]. For a given vehicular application, the system continuously monitors integrity risk, defined as the percentage of time it fails to raise an alert when the true position error exceeds the alert limit, by evaluating positioning accuracy in real time. If the system is unable to provide a reliable bound on the true position error over a predefined time window (i.e., a continuity breach), the integrity monitoring framework transitions to a fallback state. In simpler terms, the positioning solutions are made unavailable rather than being used for motion planning and control, which could be hazardous.
The research on integrity monitoring emerged alongside Global Positioning Satellite (GPS) technology in the field of aviation back in the 1980s, to provide a reliable navigation aid for aircraft. The Receiver Autonomous Integrity Monitoring (RAIM) [9] algorithm was devised for single-fault detection using redundant pseudorange measurements from satellites of a single constellation, originally GPS. Over the years, other versions of RAIM, such as advanced RAIM (ARAIM), solution separation RAIM were developed for multi-constellation and multi-fault detection and isolation. However, RAIM cannot be directly applied to CAV applications because of its foundational assumptions, threat models, and operational constraints that are more suited for aviation. Unlike aviation, automotive navigation faces challenges such as multipath, non-line-of-sight, etc. Moreover, the positioning requirements in aviation are generally less stringent than those of safety-critical CAV applications. With the advent of multi-modal sensors and V2X information sharing, it is increasingly important to develop integrity monitoring frameworks that account for multi-sensor fusion and the performance characteristics of V2X communication.
However, in mixed-traffic scenarios, there is significant variation in vehicle types, sensor configurations, levels of cooperation, and positioning techniques. It is therefore unlikely that two vehicles equipped with the same sensors and computational hardware will exhibit identical navigation performance [10]. Hence, it is essential to study and analyze the integrity of positioning systems under these varying conditions by surveying existing work and identifying key insights that can inform the development of integrity monitoring frameworks tailored to CAV-centric applications. The main contributions of this paper are:
  1. 1.
    Existing integrity monitoring work and key automotive safety standards are surveyed, covering both standalone and cooperative approaches.
  2. 2.
    The main limitations in current IM research are identified and discussed, particularly in multisensor fusion, mixed traffic scenarios, and the lack of cooperative IM solutions.
  3. 3.
    Emphasis is placed on developing cooperative integrity monitoring frameworks that leverage shared perception and V2X data in mixed-traffic environments.
  4. 4.
    The paper provides a detailed discussion linking vehicular positioning uncertainties to Required Navigation Performance (RNP) metrics, and outlines key IM techniques including RAIM, Kalman filter residuals, model-based, and set-theoretic approaches.
  5. 5.
    Existing V2X and cooperative datasets are reviewed to highlight their relevance and potential for benchmarking IM algorithms in various CAV applications.
  6. 6.
    Finally, a clear set of research priorities and future directions is established for integrity monitoring research in the context of CAVs.
@@
Although mixed traffic in CAV applications includes other road users such as vulnerable road users (VRUs), this paper focuses exclusively on vehicle positioning. The different sections of this paper are organized as follows. Section II discusses various uncertainties considered in past CAV research, i.e. model uncertainty, position uncertainty, etc. Integrity monitoring methods for vehicles via standalone and cooperative operations are presented in Section III. Section IV discusses key automotive safety standards in practice. Emerging V2X datasets are discussed in Section V. Section VI explores existing research gaps and highlights potential opportunities for future work. The paper concludes with a comprehensive conclusion, synthesizing the key findings and implications of the research in Section VII.
@@ II Background: Consideration of Positioning Uncertainties in CAV Applications
Figure 1: Fundamental operations of vehicle platooning, showcasing research opportunities and challenges in protocol design, trajectory planning, and vehicle controls.
@@
Figure 2: Identified research gaps in Connected and Autonomous Vehicle (CAV) applications. Gaps are marked in brown, while the corresponding research domains are highlighted in green.
@@ III Positioning Integrity Monitoring
The integrity monitoring (IM) framework consists of a set of requirements for the navigation system that must be met to ensure safe operation. These requirements are subject to the positioning system used by the vehicle and the application it intends to perform. These requirements are referred to as Required Navigation Performance (RNP). The performance of the vehicle positioning and navigation system is quantified using four properties defined by the RNP framework - Accuracy, Availability, Continuity, and Integrity risk. These properties are quantified by four parameters [29, 30]:
@@
Figure 3: Stanford-ESA integrity diagram to monitor positioning integrity. The location of each (PL, PE) pair on the chart indicates whether the system is in nominal operation, unavailable, misleading information, or a hazardously misleading information zone.
@@
The Stanford-ESA (Error Separation Algorithm) integrity diagram [30] is a graphical tool used to illustrate and assess the integrity of GNSS-based navigation systems, particularly in aviation. Although this plot was developed for GNSS positioning systems, it can be adapted and used for other navigation solutions. The PL estimated from standard deviations in position estimates and multipliers derived from desired confidence levels is plotted against the true position error over time across different scenarios. Different operating zones in the Stanford plot are discussed in detail in [29] and [8]. Integrity risk is determined by the percentage of pairs (PL, PE) in the Hazardously Misleading Information (HMI) region without raising an alert, thus compromising vehicle safety. The different zones of the integrity diagrams, as shown in fig3, are discussed as follows:
  1. 1.
    Zone I - Nominal operation: Both the protection level and the positioning error remain well within the alert limit, with the system conservatively estimating its error (i.e., PL $>$ PE). As a result, integrity is maintained, ensuring safe operation without the need for alerts.
  2. 2.
    Zone II - System Unavailable: When the Protection Level (PL) exceeds the Alert Limit (AL), the uncertainty in the positioning estimate is too high for the application’s safety requirement. Even if the actual Position Error (PE) is below AL, the system can’t guarantee it — so it declares the service unavailable rather than risking unsafe operation. For example, GPS-only positioning in urban environments can have errors of 5–10 meters due to multipath and signal blockage, making it unreliable for lane-centering control in automated vehicles, which typically requires errors within 1 meter.
  3. 3.
    Zone III - Misleading information: When AL $>$ PE $>$ PL, the system underestimates the actual position error, leading to overconfidence. For example, in infrastructure LiDAR-based vehicle tracking, the lateral protection level (PL) might be estimated at 0.4 meters under normal conditions. However, due to partial occlusion or a sparse point cloud — such as when a vehicle is partially blocked by another object — the actual lateral position error (PE) may jump to 1.0 meter. If the alert limit (AL) is 1.5 meters (i.e., half the lane width), the system still considers the output safe. Yet, this unnoticed degradation poses a risk, as the error is significantly higher than the system claims, reducing the effective safety margin and leading to misleading information.
  4. 4.
    Zone IV - Unavailable and misleading: Under this condition, the system not only fails to meet the required accuracy (PL $>$ AL), but also underestimates how bad the error is (PE $>$ PL). This is a critical situation as the system is both unavailable and misleading.
  5. 5.
    Zone V - Hazardously misleading information: Hazardously Misleading Information (HMI) occurs when the system underestimates its actual positioning error, resulting in false confidence. For example, in lane-level positioning, the vehicle may have already drifted out of its lane (PE $>$ AL), but the system reports a protection level (PL) well within the alert limit (PL $<$ AL). Since no alert is triggered, the system continues to operate under the assumption that the position is safe. This undetected failure poses a serious risk in safety-critical applications.
Fig4 illustrates how a vehicle’s positioning degrades from normal operation into an integrity-risk state. This shift can result from erroneous measurements—such as GNSS signal blockage or multipath errors—that undermine the position-estimation process. For this reason, GNSS is commonly fused with onboard sensors or supplemented by cooperative V2X information.
Figure 4: Vehicle’s positioning performance transitioning from nominal operation to misleading information and finally to hazardously misleading information. During the misleading stage, the system exhibits false confidence despite increased error. In the hazardously misleading stage, it fails to trigger an alert as the protection level remains below the alert limit.
III-A Overview of Key Integrity Monitoring Techniques
Integrity monitoring ensures that the estimated position from a localization system remains reliable, particularly in safety-critical applications such as autonomous driving and aviation. The goal is to detect faults or inconsistencies in sensor data or system behavior and to verify whether the estimated state stays within acceptable limits. This section outlines four widely used integrity monitoring techniques: Receiver Autonomous Integrity Monitoring (RAIM), originally developed for GNSS; Kalman Filter-based methods, which extend RAIM to dynamic systems; model or coherence-based approaches that check consistency across independent sensor sources; and set-theoretic methods, such as interval analysis, which rely on bounded-error assumptions rather than probabilistic models (. Each method is suited to different use cases and offers specific strengths depending on system design and sensor characteristics.
  1. 1.
    Receiver Autonomous Integrity Monitoring (RAIM):
    RAIM checks the consistency of pseudorange measurements received from multiple satellites by evaluating the residuals between observed and expected values. Classical RAIM, developed in the 1980s, is designed to detect a single satellite fault within a single GNSS constellation, originally GPS. In contrast, Advanced RAIM (ARAIM) extends this capability to handle multiple faults and supports multi-constellation setups. For a basic 3D position fix, signals from at least four satellites are required, while implementing RAIM requires at least five. The expected pseudoranges are computed based on the receiver’s estimated position and satellite geometry; any significant discrepancy between these and the actual measurements results in large residuals, signaling a potential fault. When such faults are detected, the system raises an integrity alert to either isolate or exclude the affected satellite’s data from the position solution.
    The least squares (LS) and weighted least squares (WLS) methods are two of the popular methods to compute residuals in pseudorange measurements and detect faults. The linearized GNSS pseudorange observation equation is formulated as:
    $y=Hx+\epsilon$ (1)
    , where y is an nx1 vector containing actual pseudorange measurements minus expected range measurements based on the location of the (n) satellites and estimated receiver position, x is an unknown 4x1 vector including receiver position and clock bias; H is an nx4 observation matrix; $\epsilon$ is an n-dimensional column vector storing measurement noises; W is the weighted matrix represented by, $\mathbf{W}=\mathrm{diag}\left(\frac{1}{\sigma_{1}^{2}},\frac{1}{\sigma_{2}^{2}},\ldots,\frac{1}{\sigma_{n}^{2}}\right)$. The estimated state vector and pseudorange residuals are computed as follows,
    $x_{wls}=(H^{T}WH)^{-1}H^{T}Wy$ (2)
    $w=y-Hx_{wls}$ (3)
    The sum of squared errors could be calculated from the residuals by,
    $SSE=w^{T}w$ (4)
    The test statistic ($T_{s}$) used for fault detection could be computed using,
    $T_{s}=\sqrt{SSE/(n-4)}$ (5)
    The test statistics follows a chi-square distribution with (n-4) degrees of freedom. Given the probability of false alarms ($P_{fa}$) the test threshold (T) can be computed using eq. 6 and eq. 7,
    $P(SSE<t)=1-P_{fa}$ (6)
    $T=\sqrt{t/(n-4)}$ (7)
    If the test statistic ($T_{s}$) is less than the threshold ($T$), no fault is detected. However, if $T_{s}>T$, a fault is present and must be identified using appropriate fault isolation techniques. Apart from least squares techniques, there exist other fault detection (FD) techniques used by other variants of RAIM algorithm. These include solution separation (SS), multi-hypothesis solution separation, chi-square method, ambiguity resolution algorithm (LAMBDA) and many more. Readers are encouraged to refer to surveys [9, 31, 32] to gather additional information on advanced RAIM (ARAIM). relative RAIM, extended RAIM, carrier-based RAIM, time RAIM, vision-aided RAIM.
  2. 2.
    Kalman Filter Residual-based IM:
    Kalman filter is a special case of least-squares estimators, specifically, an optimal, recursive, weighted least squares estimator designed for dynamic systems under Gaussian noise assumptions. At each time instance $t_{k}$, the innovation or residual $r_{k}$ is calculated as the difference between the actual measurement $z_{k}$ and the predicted measurement $\hat{z}_{k}=H\hat{x}_{k|k-1}$,
    $r_{k}=z_{k}-H\hat{x}_{k|k-1}$ (8)
    , where $\hat{x}_{k|k-1}$ is the predicted state and H is the observation matrix. The covariance of the residual is computed as,
    $S_{k}=H_{k}P_{k|k-1}H_{k}^{T}+R_{k}$ (9)
    with $P_{k|k-1}$ is the predicted error covariance and $R_{k}$ is the measurement error covariance. A chi-squared test statistic is defined as,
    $T_{k}=r^{T}_{k}S^{-1}_{k}r_{k}$ (10)
    Similar to the least squares method used in RAIM, a fault is declared if the test statistics $T_{k}$ exceed the test threshold corresponding to a given false alarm probability. While traditional RAIM methods rely on least squares estimation, Kalman filter-based residuals have been explored as an alternative for fault detection in multi-sensor fusion systems combining GNSS with sensors such as cameras and LiDAR [26]. For non-linear systems, Variants of the Kalman filter such as Extended KF, Unscented KF could be used in [33, 34].
    Figure 5: Schematic overview of four integrity-monitoring approaches: (a) Receiver Autonomous Integrity Monitoring (RAIM) evaluates pseudorange residuals to detect satellite faults; (b) Kalman filter residual–based IM calculates the innovation for filter consistency; (c) Model-based IM performs cross-consistency checks by fitting environmental features (map waypoints, GPS traces, camera-detected lane markings) across multiple sensors; and (d) Set-theoretic IM uses interval-based methods to compute feasible solution sets that provide a guaranteed bound for the true vehicle position.
  3. 3.
    Model & Coherence based IM:
    Localization errors in perception sensors can arise not only at the system or application level but also at the signal or data level [35]. Therefore, it is important to cross-check data consistency at the sensor source to prevent the propagation of errors into downstream modules. The key idea behind model consistency or coherence estimation across multiple sensor data streams is to assess the similarity in their perception of external environmental features. These features may include lane markings, curbs, landmarks, or other identifiable markers. In [24, 36], a cross-consistency check was performed to evaluate Goodness-of-Fit (GoF) among data sources.
    Given $N$ spatial sensor sources denoted as $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{N}\}$, each providing 2D data points for environmental features (curb, lanes, etc.) in a common reference frame (e.g., ego-frame). $\mathbf{D}_{i}=\{(x_{i1},y_{i1}),\ldots,(x_{iM_{i}},y_{iM_{i}})\}\subset\mathbb{R}^{2}$ represents spatial data from the sensor $S_{i}$. The common feature (e.g., lane marking) is fitted with a second-degree polynomial model:
    $p_{i}(x)=a_{i}x^{2}+b_{i}x+c_{i}$ (11)
    , where $p_{i}$ represents the model and ${a_{i},b_{i},c_{i}}$ represent the polynomial coefficients estimated via robust regression. To evaluate the model $p_{i}$ on data from sensor $S_{j}$, a cross-model polynomial is defined:
    $p_{ij}(x)=a_{i}x^{2}+b_{i}x+c_{j}$ (12)
    , where $c_{j}$ is taken from $p_{j}$ to align the offset to match the data from $S_{j}$. The curvature at point $x$ for model $p_{ij}$ is defined as:
    $\kappa_{ij}(x)=\frac{2a_{i}}{\left(4a_{i}^{2}x^{2}+4a_{i}b_{i}x+b_{i}^{2}+1\right)^{3/2}}$ (13)
    The computed curvature is used to weight the residuals. The weighted residual error for model $p_{ij}$ on data $\mathbf{D}_{j}$ is computed as:
    $e_{ij}=\frac{1}{M_{j}}\sum_{k=1}^{M_{j}}\kappa_{ij}(x_{jk})\cdot\left|y_{jk}-p_{ij}(x_{jk})\right|$ (14)
    The error $e_{ij}$ is the GoF marker and quantifies the disagreement of sensor $i$’s model with data from sensor $j$. All pairwise cross-consistency values are collected in matrix $E\in\mathbb{R}^{N\times N}$:
    $E=[e_{ij}],\quad i,j=1,\ldots,N$
    Integrity markers $W_{i}$ for each sensor $S_{i}$ is defined as:
    $W_{i}=1-\frac{\sum_{j=1}^{N}e_{ij}}{\sum_{i=1}^{N}\sum_{j=1}^{N}e_{ij}},$ (15)
    , such that $\sum_{i=1}^{N}W_{i}=1$. A greater agreement with other sensor sources leads to a higher value of $W_{i}$. Any discrepancy or fault in sensor information would lead to a higher residual and a lower integrity score.
  4. 4.
    Set Theoretic IM:
    The aforementioned IM techniques belong to fault detection (FD) methods, which rely on probabilistic models to compute the protection level (PL). However, the bounded error (BE) approaches provide a deterministic solution on the error bound and thus can work with unknown or non-gaussian noise distributions [37, 38]. It also guarantees the containment of measurements within a bound often represented using sets, which is critical in unpredictable and safety-critical scenarios involving autonomous driving [39].
    Bounded error approaches include interval analysis, zonotope estimation, set-membership estimation, set-inversion via interval analysis (SIVIA), etc. Among these, interval analysis is a set-theoretic technique where uncertainties are modeled as closed intervals rather than probabilistic distributions [29]. Each scalar variable $x$ is represented as an interval, $[x]=[x_{min},x_{max}]$, and a vector of variables $\mathbf{x}\in\mathbb{R}^{n}$ is described using a n-dimensional box, which is the Cartesian product of scalar intervals:
    $[\mathbf{x}]=[x_{1}]\times[x_{2}]\times\dots\times[x_{n}].$ (16)
    A scalar function $f(x)$ can be extended to an interval function $[f]([x])$, which evaluates as:
    $[f]([x])=\{f(x)\mid x\in[x]\}.$ (17)
    An inclusion function is one that always contains the true image of the function, meaning:
    $f([x])\subseteq[f]([x]).$ (18)
    In most practical cases, inclusion functions over-approximate the actual range due to variable dependencies or nonlinearity, but they ensure that the true value is always enclosed. In the context of estimation, interval analysis can be used to solve constraint satisfaction problems (CSPs) where the objective is to find the smallest box $[x]$ such that all constraint functions $f_{i}(x)\in[z_{i}]$ are satisfied. When some measurements are suspected to be outliers, a $q$-relaxed CSP can be formulated, which allows up to $q$ constraints to be violated. The integrity risk associated with this relaxation can be expressed as:
    $R=1-\sum_{k=m-q}^{m}\binom{m}{k}p^{k}(1-p)^{m-k},$ (19)
    where $m$ is the total number of constraints, $q$ is the maximum number of tolerated outliers, and $p$ is the probability that a measurement lies within its bounded interval. The method is particularly useful for integrity monitoring in localization tasks, as it provides hard guarantees even under bounded noise and measurement faults.
III-B In-vehicle (Standalone) Integrity Monitoring